# MixGRPO: Unlocking Flow-based GRPO Efficiency with Mixed ODE-SDE

### v2

@@ 5 Conclusion
Although GRPO (Shao et al., 2024) has seen significant success in the language modality, it is still in the early stages of progress in vision (Tong et al., 2025; Xue et al., 2025; Liu et al., 2025a). Existing flow-based GRPO faces challenges such as low sampling efficiency and slow training. To address these issues, we proposed MixGRPO, a novel training framework that combines SDE and ODE sampling. This hybrid approach allows for focused optimization on the SDE sampling flow component, reducing complexity while ensuring accurate reward computation. Inspired by the decay factor in reinforcement learning (Hu et al., 2022b), we introduce a sliding window strategy for scheduling the optimized denoising steps. Experimental results confirm the effectiveness of our approach in both single-reward and multi-reward settings. Additionally, MixGRPO decouples the denoising stages for optimization and reward computation, enabling acceleration of the latter with high-order solvers. We further propose MixGRPO-Flash, which balances overhead and performance. We hope MixGRPO will inspire deeper research into post-training for image generation, contributing to the advancement of Artificial General Intelligence (AGI).
6 Acknowledgement
The authors thank Yiming Cheng (M.Sc., Tsinghua University, 13306339695@163.com) for his valuable feedback on the theoretical derivations and mathematical expressions.
Appendix A Proof of convergence for Mixed ODE-SDE sampling
To prove that the mixed ODE-SDE sampling method in Eq. (4) has the same convergence as Eq. (2), which uses only ODE sampling, referencing Song et al. (2020b), we approach this from the perspective of distribution evolution, where the distribution at each time step, e.g., $\frac{\partial{q_{t}(\mathbf{x})}}{\partial{t}}$ must be the same. Let the interval for SDE be defined as $S=\left[t_{l},t_{r}\right)\in\left[0,1\right)$. Along the denoising direction, when the same initial Gaussian noise distribution $q_{0}(\mathbf{x}_{0})$ is given, the probability distribution evolution in the ODE interval preceding the SDE is completely identical. The key point is whether the distribution evolution of the SDE within the interval $S$ is completely equivalent to that of the ODE. If they are equivalent, then the ODE interval following the SDE will naturally be equivalent to using only ODE sampling. Next, we will provide a detailed proof for this key point.
@@ Appendix A Proof of convergence for Mixed ODE-SDE sampling
The marginal probability density $q_{t}(\mathbf{x}_{t})$ evolves according to Kolmogorov’s equation (Fokker-Planck equation) (Øksendal, 2003)
$\frac{\partial q_{t}(\mathbf{x})}{\partial t}=-\nabla_{\mathbf{x}}\cdot\big[\big(f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)q_{t}(\mathbf{x})\big]+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$ (13)
@@
$\frac{\partial q_{t}(\mathbf{x})}{\partial t}$ $=-\nabla_{\mathbf{x}}\cdot\big[f(\mathbf{x},t)q_{t}(\mathbf{x})-g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})\big]+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$
@@
$=-\nabla_{\mathbf{x}}\cdot\big[\underbrace{\big(f(\mathbf{x},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)}_{f_{\text{ODE}}(\mathbf{x},t)}q_{t}(\mathbf{x})\big].$ (14)
The Eq. (A) is indeed the Fokker-Planck equation of the ODE Eq. (2). Therefore, within the interval $S$, the distribution evolution of SDE and ODE sampling is consistent.
Appendix B Hybrid Inference for Solving Reward Hacking
To address the reward hacking issue, we employ hybrid inference to sample with both the raw and post-trained models, and introduce the hybrid percent $p_{\text{mix}}$. This means that the initial $p_{\text{mix}}T$ denoising steps are sampled by the model trained with GRPO, while the remaining denoising process is finished by the original FLUX model (Black et al., 2023). Table 10 and Figure 4 respectively illustrate the changes in performance and images as $p_{\text{mix}}$ increases under the multi-rewards training scenario. The experimental results demonstrate that $p_{\text{mix}}=80\%$ is an optimal empirical value that effectively mitigates hacking while maximizing alignment with human preferences.
Table 10: Comparison with different hybrid inference percentages
$p_{\text{mix}}$ HPS-v2.1 Pick Score ImageReward Unified Reward
0% 0.313 0.226 1.089 3.369
20% 0.342 0.233 1.372 3.386
40% 0.356 0.235 1.539 3.395
60% 0.362 0.236 1.598 3.407
80% 0.366 0.238 1.610 3.411
100% 0.369 0.238 1.607 3.378
Figure 4: Qualitative comparison with different hybrid inference percentages
Appendix C DPM-Solver++ for Recitified Flow
@@
The DPM-Solver++ algorithm (Lu et al., 2022b) is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model (Rombach et al., 2022), where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t_{i}}$, the time condition $t_{i}$ and the text condition $c$. According to the definition of Rectified Flow (RF) (Liu et al., 2022), there is the following transfer equation:
@@
According to the theory of stochastic interpolation (Albergo et al., 2023), RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t_{i}}$:
@@
Taking the multistep second-order DPMSolver++ as an example (see Algorithm 2 in  (Lu et al., 2022b)), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{i}$:
@@
Based on the exact discretization formula for the probability flow ODE proposed in DPM-Solver++ (Eq. (9) in (Lu et al., 2022b)), we can derive the final transfer equation:
@@
Appendix D MixGRPO Algorithm
Taking the progressive movement strategy of the sliding window as an example, the MixGRPO Algorithm 1 is as follows. MixGRPO first sets the left boundary $l$ of the sliding window $W(l)$ at $t=0$ of the reverse process, which corresponds to the first sampling time-step. A GRPO iteration consists of two phases: sampling and optimization.
In the sampling phase, mixed ODE-SDE sampling is performed for the $N$ samples within the group, using SDE within the window and ODE outside. Then, the reward and advantage for the $N$ samples are calculated. For the multi-reward strategy, the advantage for each reward model is computed separately and then aggregated to obtain the final advantage. In the optimization phase, the GRPO loss (Shao et al., 2024) is computed and optimized only at the sampling timesteps within the SDE window $W(l)$.
After each GRPO iteration, the current training steps are checked against the conditions for moving the window, determining whether to shift the window by the window stride $s$.
Algorithm 1 MixGRPO Training Process
3:Init left boundary of $W(l)$: $l\leftarrow 0$
4:for training iteration $m=1$ to $M$ do
5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
9:   for generate $i$-th image from $i=1$ to N do
10:     for sampling timestep $t=0$ to T1T-1 do $\triangleright$ $\pi_{\theta_{\text{old}}}$ mixed sampling loop
11:      if $t\in W(l)$ then
12:        Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
13:      else
14:        Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
15:      end if
16:     end for
17:   end for
18:   for $i$-th image from $i=1$ to N do
19:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
20:   end for
21:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
22:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
23:   end for
24:  end for
26:   $l\leftarrow\min(l+s,~T-w)$
27:  end if
28:end for
Appendix E MixGRPO-Flash Algorithm
@@
16:        Use higher-order ODE sampling to get $\mathbf{x}^{i}_{t+1}$
@@
21:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
@@
23:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
24:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
@@
Appendix F More visualized results
Figure 5: Qualitative comparison with different training-time sampling steps. The performance of MixGRPO does not significantly decrease with the reduction in overhead. The Frozen strategy means that optimization is only employed at the initial denoising steps.

### v3

@@ 4.4.2 High Order ODE Solver
5 Limitation
A key limitation of GRPO is that its performance ceiling is inherently tied to the capability of the reward model. Specifically, when a reward model is not sufficiently powerful to provide a comprehensive assessment of image quality, the GRPO algorithm is prone to reward hacking, particularly in the later stages of training. As confirmed by several studies [39, 38, 23], reward hacking is fundamentally an issue stemming from the limitations of the reward model and is less attributable to the RL algorithm itself. In this context, the primary focus of MixGRPO is not to overcome the reward model’s limitations, but rather to explore how to achieve faster convergence and superior performance under the guidance of existing, imperfect reward models. Looking ahead, while the ultimate performance is contingent upon the reward model’s quality, we plan to explore and train more powerful reward models as a direction for future work.
@@ 6 Conclusion
In this work, we have presented MixGRPO, a novel hybrid ODE-SDE framework for improving GRPO training efficiency and performance. We have proposed a strategy to confine optimization to a dynamic stochastic interval managed by a sliding window. This guides the optimization process from broad exploration to fine-grained refinement, thus enhancing performance. Experiments demonstrate that MixGRPO has achieved superior performance in both single-reward and multi-reward settings while substantially reducing overhead. Furthermore, we have presented MixGRPO-Flash, a variant offering a flexible trade-off between performance and computational cost.
Supplementary Material
7 Proof of Convergence for Mixed ODE-SDE Sampling
To prove that the mixed ODE-SDE sampling method in Eq. (7) has the same convergence as Eq. (2), which uses only ODE sampling, referencing [34], we approach this from the perspective of distribution evolution, where the distribution at each time step, e.g., $\frac{\partial{q_{t}(\mathbf{x})}}{\partial{t}}$ must be the same. Let the interval for SDE be defined as $S=\left[t_{l},t_{r}\right)\in\left[0,1\right)$. Along the denoising direction, when the same initial Gaussian noise distribution $q_{0}(\mathbf{x}_{0})$ is given, the probability distribution evolution in the ODE interval preceding the SDE is completely identical. The key point is whether the distribution evolution of the SDE within the interval $S$ is completely equivalent to that of the ODE. If they are equivalent, then the ODE interval following the SDE will naturally be equivalent to using only ODE sampling. Next, we will provide a detailed proof for this key point.
@@ 7 Proof of Convergence for Mixed ODE-SDE Sampling
The marginal probability density $q_{t}(\mathbf{x}_{t})$ evolves according to Kolmogorov’s equation (Fokker-Planck equation) [24]
$\frac{\partial q_{t}(\mathbf{x})}{\partial t}=$ $-\nabla_{\mathbf{x}}\cdot\big[\big(f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)q_{t}(\mathbf{x})\big]$ (13)
$+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$
@@
$\frac{\partial q_{t}(\mathbf{x})}{\partial t}$ $=-\nabla_{\mathbf{x}}\cdot\big[f(\mathbf{x},t)q_{t}(\mathbf{x})-g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})\big]$ (14)
$\qquad~+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$
@@
$=-\nabla_{\mathbf{x}}\!\cdot\!\big[\!\underbrace{\big(f(\mathbf{x},t)\!-\!\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)}_{f_{\text{ODE}}(\mathbf{x},t)}q_{t}(\mathbf{x})\!\big].$
The Eq. (14) is indeed the Fokker-Planck equation of the ODE Eq. (2). Therefore, within the interval $S$, the distribution evolution of SDE and ODE sampling is consistent.
8 DPM-Solver++ for Recitified Flow
@@
The DPM-Solver++ algorithm [22] is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model [29], where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t_{i}}$, the time condition $t_{i}$ and the text condition $c$. According to the definition of Rectified Flow (RF) [19], there is the following transfer equation:
@@
According to the theory of stochastic interpolation [1], RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t_{i}}$:
@@
Taking the multistep second-order DPMSolver++ as an example (see Algorithm 2 in  [22]), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{i}$:
@@
Based on the exact discretization formula for the probability flow ODE proposed in DPM-Solver++ (Eq. (9) in [22]), we can derive the final transfer equation:
@@
9 MixGRPO-Flash Algorithm
@@
16:         Use DPM-Solver++ sampling to get $\mathbf{x}^{i}_{t+1}$
@@
21:     Calculate advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
@@
23:    for optimization timestep $t\in W(l)$ do
24:     Update policy model: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
@@
10 Hybrid Inference for Solving Reward Hacking
As discussed in Section 5, reward hacking stems from the limited evaluation capabilities of the reward model. To address reward hacking and improve visualization, we employ the hybrid inference strategy from [43], which uses the post-trained model for low-SNR (signal-to-noise ratio) steps and the original model for high-SNR steps during inference-time sampling. In our experiments, we also applied hybrid inference to the other baseline models to ensure a fair and consistent comparison.
We employ hybrid inference and introduce the hybrid percent $p_{\text{mix}}$. This means that the initial, high-SNR $p_{\text{mix}}T$ denoising steps, are handled by the model trained with GRPO, while the remaining denoising process is finished by the original model. Table 10 and Figure 5 respectively illustrate the changes in performance and images as $p_{\text{mix}}$ increases under the multi-rewards training scenario. The experimental results demonstrate that $p_{\text{mix}}=80\%$ is an optimal empirical value that effectively mitigates hacking while maximizing alignment with human preferences.
Table 10: Comparison with different hybrid inference percentages
$p_{\text{mix}}$ HPS-v2.1 Pick Score ImageReward Unified Reward
0% 0.313 0.226 1.089 3.369
20% 0.342 0.233 1.372 3.386
40% 0.356 0.235 1.539 3.395
60% 0.362 0.236 1.598 3.407
80% 0.366 0.238 1.610 3.411
100% 0.369 0.238 1.607 3.378
Figure 5: Qualitative comparison with different hybrid inference percentages. We found 80% to be the optimal value, as it maximizes image quality without causing over-optimization.
11 Cross-Dataset Experiments
To investigate the robustness and parameter sensitivity of the sliding window strategy in MixGRPO, we conducted a series of cross-dataset ablation studies. We established two reciprocal settings to evaluate both in-domain (ID) and out-of-domain (OOD) performance. In cross-dataset experiment 1, the model was trained on the HPDv2 [40] dataset and evaluated on the test sets of HPDv2 (ID) and Pick-a-Pic v1 (OOD). In cross-dataset experiment 1, conversely, the model was trained on Pick-a-Pic v1 [12] and evaluated on the test sets of Pick-a-Pic v1 (ID) and HPDv2 (OOD). Within these settings, we ablated the key parameters of MixGRPO: the moving strategy, shift interval $\tau$, window size $w$ and window stride $s$. The results are presented in the Tables 11121314.
As shown in Table 11, Progressive-Decay(Exp) and Progressive-Constant emerge as the top-performing moving strategies. While Progressive-Decay(Exp) consistently achieves the highest ImageReward [41] score in in-domain evaluations, Progressive-Constant demonstrates superior overall performance. Notably, the Progressive-Constant strategy consistently outperforms all others in out-of-domain tests, demonstrating its strong robustness and generalization capabilities.
The results concerning the shift interval ($\tau$), which dictates the frequency of window movement in training iteration steps, are summarized in Table 12. Overall, the optimal performance is consistently observed at $\tau=25$. Specifically, as $\tau$ is incrementally increased from $15$ to $25$, performance metrics exhibit a gradual rise across both in-domain and out-of-domain test sets. However, a crucial observation is the sharp decline in performance when the interval is extended to $\tau=30$, with results even falling below the baseline achieved at $\tau=15$. This trend suggests that a moderate reduction in the sliding window’s movement speed (from $\tau=15$ to $\tau=25$) allows the temporal behavior within the window to be sufficiently optimized. Conversely, overly slow window movement (exceeding $\tau=25$) is prone to over-optimization at certain timesteps, causing the model’s distribution to diverge significantly from the target preferences of the reward models. The cross-dataset experiments further confirm the universality of $\tau=25$ as the optimal setting.
The comparative results for the window size ($w$) are presented in Table 13. It is evident that both $w=4$ and $w=6$ represent optimal settings. Specifically, $w=6$ slightly surpasses $w=4$ only on the in-domain test set of Cross-dataset Experiment 1, and in other experiments, it only holds a marginal advantage in the HPS-v2.1 [40] score. It is crucial to note that the window size $w$ corresponds to the number of denoising timesteps that must be optimized during each training iteration, which, in turn, is linearly correlated with the overall optimization overhead of the Reinforcement Learning (RL) process. Consequently, $w=4$ is identified as the best compromise, offering an optimal trade-off between performance and computational efficiency. Regarding parameter sensitivity, we observe that the change in performance metrics—across both in-domain and out-of-domain tests—is relatively minimal when $w$ varies between $2$ and $6$. This stability suggests that the model exhibits robust performance with respect to the window size setting. As long as the window size remains within a reasonable range (e.g., $2\leq w\leq 6$), the model is capable of effective learning and maintaining stable performance, thereby reducing necessity for meticulous hyperparameter tuning.
The results concerning the stride ($s$), which defines the step size for each sliding window movement, are presented in Table 14. Considering all experimental metrics across the different datasets, a stride of $s=1$ is identified as the optimal setting overall. It is important to note that the optimal window size of $w=4$ was used in this experiment. With $s=1$, the last three timesteps within the window are repeatedly optimized in the subsequent training cycle. Crucially, the experimental results indicate that this repetition leads to performance improvement. This is likely due to the nature of the high-SNR denoising process, where the initial timesteps (e.g., the first timestep in the window) are subject to a greater extent of GRPO optimization, despite the window size being set at $w=4$. Conversely, the behavior learned during the subsequent three timesteps remains relatively under-optimized (underfitted) and thus significantly benefits from this repeated optimization. Furthermore, we observe that the highest in-domain HPS-v2.1 [40] score is achieved when $s=3$. This localized optimal result may be attributed to the HPS-v2.1 [40] reward model’s preference alignment being relatively easier to optimize, suggesting that repeated optimization (i.e., smaller strides) could potentially induce over-optimization in this specific metric.In summary, $s=1$ proves to be the most robust and generalizable choice across the varying experimental conditions.
Table 11: Comparison for different moving strategies in cross-dataset experiments. The optimal setting is highlighted in green.
Strategy Interval Schedule Cross-dataset Experiment 1 Cross-dataset Experiment 2
In Domain Dataset (HPD v2) Out-of-Domain Dataset (Pick-a-Pi v1) In Domain Dataset (Pick-a-Pi v1) Out-of-Domain Dataset (HPD v2)
HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward
Frozen / 0.354 0.234 1.580 0.352 0.226 1.539 0.346 0.230 1.601 0.351 0.231 1.587
Random Constant 0.365 0.237 1.513 0.361 0.230 1.512 0.349 0.227 1.524 0.343 0.225 1.530
Progressive Decay(Linear) 0.365 0.235 1.566 0.363 0.229 1.572 0.363 0.231 1.614 0.355 0.230 1.597
Decay(Exp) 0.360 0.239 1.632 0.364 0.232 1.612 0.361 0.234 1.628 0.358 0.230 1.584
Constant 0.367 0.237 1.629 0.366 0.234 1.622 0.365 0.238 1.618 0.359 0.232 1.601
Table 12: Comparison for different shift intervals $\tau$ in cross-dataset experiments. The optimal setting is highlighted in green.
$\tau$ Cross-dataset Experiment 1 Cross-dataset Experiment 2
In Domain Dataset (HPD v2) Out-of-Domain Dataset (Pick-a-Pi v1) In Domain Dataset (Pick-a-Pi v1) Out-of-Domain Dataset (HPD v2)
HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward
15 0.366 0.237 1.509 0.359 0.227 1.470 0.358 0.234 1.610 0.353 0.228 1.542
20 0.366 0.238 1.610 0.360 0.228 1.619 0.361 0.237 1.615 0.357 0.233 1.588
25 0.367 0.237 1.629 0.366 0.234 1.623 0.365 0.238 1.618 0.359 0.232 1.601
30 0.350 0.229 1.589 0.348 0.221 1.585 0.355 0.234 1.609 0.351 0.229 1.509
Table 13: Comparison for different window sizes $w$ in cross-dataset experiments. The optimal setting is highlighted in green.
$w$ Cross-dataset Experiment 1 Cross-dataset Experiment 2
In Domain Dataset (HPD v2) Out-of-Domain Dataset (Pick-a-Pi v1) In Domain Dataset (Pick-a-Pi v1) Out-of-Domain Dataset (HPD v2)
HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward
1 0.359 0.232 1.571 0.349 0.223 1.445 0.349 0.229 1.553 0.346 0.221 1.565
2 0.366 0.235 1.618 0.363 0.230 1.614 0.363 0.234 1.597 0.359 0.230 1.585
4 0.367 0.237 1.629 0.366 0.234 1.623 0.365 0.238 1.618 0.359 0.232 1.601
6 0.370 0.238 1.624 0.364 0.232 1.620 0.366 0.237 1.608 0.362 0.232 1.594
Table 14: Comparison for different window strides $s$ in cross-dataset experiments. The optimal setting is highlighted in green.
$s$ Cross-dataset Experiment 1 Cross-dataset Experiment 2
In Domain Dataset (HPD v2) Out-of-Domain Dataset (Pick-a-Pi v1) In Domain Dataset (Pick-a-Pi v1) Out-of-Domain Dataset (HPD v2)
HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward
1 0.367 0.237 1.629 0.366 0.234 1.623 0.364 0.238 1.618 0.359 0.232 1.601
2 0.357 0.236 1.575 0.357 0.233 1.587 0.363 0.234 1.574 0.354 0.231 1.591
3 0.370 0.236 1.578 0.364 0.230 1.579 0.368 0.237 1.586 0.358 0.228 1.585
4 0.368 0.238 1.575 0.349 0.224 1.573 0.366 0.231 1.566 0.357 0.225 1.568
12 More Visualized Results